OpenStudy (marcelie):

Help me !!!! For the following exercises solve the equation below and express the answer using set notation. l 6-x l = 5 But the answer is 1, 1/2 , -2. How did they get these answers ?

OpenStudy (marcelie):

This is what i did : l 6-x l = 5 -x +6 = 5 x = 1 ? please tell me if i am correct or wrong ?

OpenStudy (sohailiftikhar):

you are correct

imqwerty (imqwerty):

yea x=1 or x=11 are the answers :) u are 100% correct :D

OpenStudy (zzr0ck3r):

I would not say you are 100% correct. You are 50% correct. Remember \(|A|=b\) means \(a=b\) and \(-a=b\) So you have \(6-x=5\) which gives \(x=1\) But we must also consider \[-(6-x)=5\] which leads to \(x=11\).

OpenStudy (sohailiftikhar):

yes

OpenStudy (zzr0ck3r):

But we are all actually wrong as it asked for the answer in set notation, so the answer is \(\{1,11\}\)

OpenStudy (marcelie):

so the interval notation is from smallest to biggest ?

imqwerty (imqwerty):

\[\left\{ 1,11 \right\}\]means that there r 2 elementary solutions 1 and 11 :)

OpenStudy (sohailiftikhar):

when ever you see the modulus you have to consider the negative and positive values means less then the positive value and greater than the negative value like that \[\left| x \right|=4\] it means -4<x>4

OpenStudy (marcelie):

oh okay . Got it !! Thank Yous !!! I have another equation bur under same category. l 4x-2 l =11 x= 13/4 and x = -9/4 ? interval notation : { -9/4 , 13/4 }

OpenStudy (anonymous):

You are correct

OpenStudy (marcelie):

for my second problem ?

OpenStudy (sohailiftikhar):

yes! :) now you got the concept buddy

OpenStudy (sohailiftikhar):

now close this question and ask new :)

OpenStudy (marcelie):

Okay. I checked my solution book and it did include the - 9/4 ?

OpenStudy (marcelie):

i mean 9/4 .

OpenStudy (sohailiftikhar):

wait in book -9/4 is not given ?

OpenStudy (marcelie):

yes , the interval notation says { 9/4 , 13/4}

OpenStudy (sohailiftikhar):

no it will be -9/4 to 13/4 not 9/4

OpenStudy (marcelie):

Okay . Thanks.

OpenStudy (sohailiftikhar):

np